A student takes an exam containing 12 multiple choice questions. Each question has 5 choices. At least 8 correct answers are required to pass. If the student guesses, what is the probability that he will pass? Round your answer to four decimal places.
Answer #1

P(correct guess) = 1/5 = 0.2

n = 12

P(X = x) = 12Cx * 0.2x * (1 - 0.2)12-x

P(X > 8) = P(X = 8) + P(X = 9) + P(X = 10) + P(X = 11) + P(X = 12)

              = 12C8 * 0.28 * 0.84 + 12C9 * 0.29 * 0.83 + 12C10 * 0.210 * 0.82 + 12C11 * 0.211 * 0.81 + 12C12 * 0.212 * 0.80

              = 0.0006